(D^2 + 4) y = 4 sec 2x csc 2x


1
Expert's answer
2021-12-14T03:24:59-0500

The corresponding homogeneous equation is


"(D^2+4)y=0"

Auxiliary equation


"r^2+4=0"

"r_1=2i, r_2=-2i"

The general solution of the homogeneous equation is


"y_h=c_1\cos(2x)+c_2\sin(2x)"

Variation of Parameters


"y'=c_1'\cos(2x)-2c_1\sin(2x)+c_2'\sin(2x)+2c_2\cos(2x)"

Let

"c_1'\cos(2x)+c_2'\sin(2x)=0"

Then


"y'=-2c_1\sin(2x)+2c_2\cos(2x)"

"y''=-2c_1'\sin(2x)-4c_1\cos(2x)"

"+2c_2'\cos(2x)-4c_2\sin(2x)"

Substitute


"-2c_1'\sin(2x)-4c_1\cos(2x)"

"+2c_2'\cos(2x)-4c_2\sin(2x)"

"+4c_1\cos(2x)+4c_2\sin(2x)"

"=4\sec(2x)\csc(2x)"

"\begin{cases}
 c_1'\cos(2x)+c_2'\sin(2x)=0\\
\\
 -2c_1'\sin(2x)+2c_2'\cos(2x)=4\sec(2x)\csc(2x)
\end{cases}"

"c_2'=-\dfrac{\cos(2x)}{\sin(2x)}c_1'"

"-c_1'\sin(2x)-\dfrac{\cos^2(2x)}{\sin(2x)}c_1'=\dfrac{2}{\sin(2x)\cos(2x)}"

"c_1'=-\dfrac{2}{\cos(2x)}"

Integrate


"c_1=-\int \dfrac{2}{\cos(2x)}dx"

"\int \dfrac{2}{\cos(2x)}dx=\int \dfrac{2\cos(2x)}{\cos^2(2x)}dx"

"=\int \dfrac{2\cos(2x)}{1-\sin^2(2x)}dx"

"=\int \dfrac{\cos(2x)}{1-\sin(2x)}dx+\int \dfrac{\cos(2x)}{1+\sin(2x)}dx"

"=-\dfrac{1}{2}\int \dfrac{d(1-\sin(2x))}{1-\sin(2x)}dx+\dfrac{1}{2}\int \dfrac{d(1+\sin(2x))}{1+\sin(2x)}dx"

"=-\dfrac{1}{2}\ln(|1-\sin(2x)|)+\dfrac{1}{2}\ln(|1+\sin(2x)|)+C_3"

"=\dfrac{1}{2}\ln\dfrac{(1+\sin(2x))^2}{1-\sin^2(2x)}+C_3"

"=\ln(1+\sin(2x))-\ln(|\cos(2x)|)+C_3"

"c_1=-\ln(1+\sin(2x))+\ln(|\cos(2x)|)-C_3"

"c_2'=\dfrac{2}{\sin(2x)}"

Integrate


"c_2=\int \dfrac{2}{\sin(2x)}dx"

"\int \dfrac{2}{\sin(2x)}dx=\int \dfrac{2\sin(2x)}{\sin^2(2x)}dx"

"=\int \dfrac{2\sin(2x)}{1-\cos^2(2x)}dx"

"=\int \dfrac{\sin(2x)}{1-\cos(2x)}dx+\int \dfrac{\sin(2x)}{1+\cos(2x)}dx"

"=\dfrac{1}{2}\int \dfrac{d(1-\cos(2x))}{1-\cos(2x)}dx-\dfrac{1}{2}\int \dfrac{d(1+\cos(2x))}{1+\cos(2x)}dx"

"=\dfrac{1}{2}\ln(|1-\cos(2x)|)-\dfrac{1}{2}\ln(|1+\cos(2x)|)+C_4"

"=-\dfrac{1}{2}\ln\dfrac{(1+\cos(2x))^2}{1-\cos^2(2x)}+C_4"

"=-\ln(1+\cos(2x))+\ln(|\sin(2x)|)+C_4"

"c_2=-\ln(1+\cos(2x))+\ln(|\sin(2x)|)+C_4"

The general solution of the homogeneous equation is


"y_h=c_5\cos(2x)+c_6\sin(2x)"


"+\bigg(-\ln(1+\sin(2x))+\ln(|\cos(2x)|)\bigg)\cos(2x)"

"+\bigg(-\ln(1+\cos(2x))+\ln(|\sin(2x)|)\bigg)\sin(2x)"